A 500 µC charge is placed at the center of a cube of sides 10 cm. Determine the electric flux through each of the sides.
_________ N·m2/C
Answer #1

flux per face = Q/6ε = 500 x 10^-6/6(8.85 x10^-12 )= 9.4 x 10^6 Nm^2 /C